OpenStudy (anonymous):

Write the following comparison as a ratio reduced to lowest terms. 10 nickels to 11 dimes

OpenStudy (anonymous):

10/11*2=10/22=5/11

jimthompson5910 (jim_thompson5910):

10 nickels to 11 dimes 10 nickels to 11*2 nickels 10 nickels to 22 nickels 10 nickels: 22 nickels 10:22 5:11 5/11

OpenStudy (anonymous):

what about 44nickels to 4 dimes

OpenStudy (anonymous):

Do the same. Multiply dimes by 2 and simplify.

jimthompson5910 (jim_thompson5910):

44nickels to 4 dimes 44nickels to 4*2 nickels 44 nickels to 8 nickels 44 nickels: 8 nickels 44:8 22:4 11:2 11/2

OpenStudy (anonymous):

how did you get 22:4

jimthompson5910 (jim_thompson5910):

I divide both components of the ratio 44:8 by 2

jimthompson5910 (jim_thompson5910):

divided*